OpenStudy (anonymous):

How do you factor 2x squared + 20x + 50?

jimthompson5910 (jim_thompson5910):

2x^2 + 20x + 50 2(x^2 + 10x + 25)

jimthompson5910 (jim_thompson5910):

now you need to factor x^2 + 10x + 25

OpenStudy (anonymous):

thanks

jimthompson5910 (jim_thompson5910):

np

OpenStudy (anonymous):

so is that now 2 (x - 5) (x-5)

jimthompson5910 (jim_thompson5910):

close, but (x-5)(x-5) = x^2 - 10x + 25

jimthompson5910 (jim_thompson5910):

we want the 10x to be positive

OpenStudy (anonymous):

so then it should be 2 (x + 5) (x+5)

jimthompson5910 (jim_thompson5910):

better

jimthompson5910 (jim_thompson5910):

the answer is either \[\Large 2(x+5)(x+5)\] or \[\Large 2(x+5)^2\]

OpenStudy (anonymous):

so the actual problem was to simplify 2x squared + 20x + 50 over x squared - 25. The answer I came up with was 2 (x+5) over (x-5) and then I turned that into 2x + 10 over x-5

OpenStudy (anonymous):

did I do anything wrong?

jimthompson5910 (jim_thompson5910):

no you did that perfectly, nice work

jimthompson5910 (jim_thompson5910):

you can leave the answer as \[\Large \frac{2(x+5)}{x-5}\] or you can distribute to get \[\Large \frac{2x+10}{x-5}\]

jimthompson5910 (jim_thompson5910):

either one works in my opinion

OpenStudy (anonymous):

and then ix 2x cubed - 128 this is factored form? 2 (x - 4) (x squared + 4x + 16)

jimthompson5910 (jim_thompson5910):

2x^3 - 128 2(x^3 - 64) 2(x-4)(x^2 + 4x + 16) ... use difference of cubes factoring formula I'm getting the same

OpenStudy (anonymous):

so I'm trying to solve for x in this equation: 10x + 3 over 5x + 2 = 1\2. I got rid of the 5x and turned the 10x into 2x and now have 2x over x + 2 = 1\2. is it all good so far?

jimthompson5910 (jim_thompson5910):

the equation is \[\Large \frac{10x+3}{5x+2} = \frac{1}{2}\] right?

OpenStudy (anonymous):

wait...it should be 5x + 6

jimthompson5910 (jim_thompson5910):

so it's \[\Large \frac{10x+3}{5x+6} = \frac{1}{2}\] ??

OpenStudy (anonymous):

correct

jimthompson5910 (jim_thompson5910):

You would cross multiply to get \[\Large \frac{10x+3}{5x+6} = \frac{1}{2}\] \[\Large 2(10x+3) = 1(5x+6)\] \[\Large 20x+6 = 5x+6\] I'll let you finish

OpenStudy (anonymous):

I got 15x = 0

jimthompson5910 (jim_thompson5910):

keep going

OpenStudy (anonymous):

then wouldn't x = 0 ?

jimthompson5910 (jim_thompson5910):

yep

OpenStudy (anonymous):

I'm trying to solve for y in this problem: 6y squared - 13y - 8 and I factored it out to (2y - 1) (3y+8) but now I'm stuck

jimthompson5910 (jim_thompson5910):

you have the signs mixed up

jimthompson5910 (jim_thompson5910):

it should factor to (2y+1)(3y-8)

OpenStudy (anonymous):

oh woops

jimthompson5910 (jim_thompson5910):

so if 6y^2-13y-8 = 0 then (2y+1)(3y-8) = 0 which allows us to say 2y+1=0 or 3y-8=0

jimthompson5910 (jim_thompson5910):

solve those two equations for y

OpenStudy (anonymous):

so y = - 1\2 and y = 8\3

jimthompson5910 (jim_thompson5910):

yep y = -1/2 or y = 8/3

OpenStudy (anonymous):

one more problem and I did part of it already. I'm solving for x and the equation is sqrt(5-x) - 3 = 0

OpenStudy (anonymous):

I multiplied both sides by the square root of 5 - x and got 5 - x = 3 times the square root of 5 - x and then tried multiplying the side with 3 by the square root of 5 - x to get rid of the square root, but something went wrong

jimthompson5910 (jim_thompson5910):

sqrt(5-x) - 3 = 0 sqrt(5-x) = 3 ( sqrt(5-x) )^2 = 3^2 5-x = 9 keep going

jimthompson5910 (jim_thompson5910):

make sure to check the answer

OpenStudy (anonymous):

cuz I ended up with 5 - x = 3 - 5 - x which is 3 = 0x

OpenStudy (anonymous):

ohhhh

OpenStudy (anonymous):

- 4 = x

jimthompson5910 (jim_thompson5910):

yep x = -4

OpenStudy (anonymous):

ok I promise this is the last one. I tried solving this with my friend but neither of us could solve it for some reason. I'm adding 3 over x - 2 + 5 over x squared + 5x + 6 and I then got 3 over x - 2 + 5 over (x + 2) (x + 3)

OpenStudy (anonymous):

is there anything else I need to do?

jimthompson5910 (jim_thompson5910):

the expression is \[\Large \frac{3}{x-2} + \frac{5}{x^2+5x+6}\] correct?

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

ok one sec

jimthompson5910 (jim_thompson5910):

hmm are you sure that it's not 3 over x+2 ? and could the +5x be -5x ?

OpenStudy (anonymous):

yes I'm sure

jimthompson5910 (jim_thompson5910):

ok just checking

OpenStudy (anonymous):

I"m just struggling so much cuz I haven't taken a math class since the spring

jimthompson5910 (jim_thompson5910):

oh so it's been nearly a year since you've taken math?

OpenStudy (anonymous):

ya...I was in the pastry program and was just focusing on that

jimthompson5910 (jim_thompson5910):

i gotcha, that's definitely a long time to be out of practice, but here's how you simplify

OpenStudy (anonymous):

but then I switched my major this quarter

jimthompson5910 (jim_thompson5910):

\[\Large \frac{3}{x-2} + \frac{5}{x^2+5x+6}\] \[\Large \frac{3}{x-2} + \frac{5}{(x+2)(x+3)}\] \[\Large \frac{3(x+2)(x+3)}{(x-2)(x+2)(x+3)} + \frac{5}{(x+2)(x+3)}\] \[\Large \frac{3(x+2)(x+3)}{(x-2)(x+2)(x+3)} + \frac{5(x-2)}{(x-2)(x+2)(x+3)}\] \[\Large \frac{3(x+2)(x+3)+5(x-2)}{(x-2)(x+2)(x+3)}\] \[\Large \frac{3(x^2+5x+6)+5(x-2)}{(x-2)(x+2)(x+3)}\] \[\Large \frac{3x^2+15x+18+5x-10}{(x-2)(x+2)(x+3)}\] \[\Large \frac{3x^2+20x+8}{(x-2)(x+2)(x+3)}\] --------------------------------------------------------------------------------- So \[\Large \frac{3}{x-2} + \frac{5}{x^2+5x+6}\] simplifies to \[\Large \frac{3x^2+20x+8}{(x-2)(x+2)(x+3)}\]

jimthompson5910 (jim_thompson5910):

being a chef wasn't that exciting anymore?

OpenStudy (anonymous):

no. It's what I decided to do when I was in 5th grade and then when it came down to choosing, it was either being a pastry chef or going into a music degree, so I tried the chef thing, but I wasn't committed enough and music is kinda my true passion

jimthompson5910 (jim_thompson5910):

i gotcha, it's probably a tough field and if you don't love it fully, then no point doing it

jimthompson5910 (jim_thompson5910):

but does that simplification make sense?

OpenStudy (anonymous):

it's really competitive and I had lots of problems with the teachers, so it wasn't like I was happy to be there every day

OpenStudy (anonymous):

it does, but it's just a really long process

jimthompson5910 (jim_thompson5910):

well maybe it was just the teachers? you never know, you might have loved the people in the actual workforce and the job itself

jimthompson5910 (jim_thompson5910):

yeah it's a lot, it would be simpler if the x-2 was x+2, but oh well

jimthompson5910 (jim_thompson5910):

or it would be simpler if the x^2 + 5x + 6 was x^2 - 5x + 6

OpenStudy (anonymous):

I got through half the program which was the basics of it...the next half is just more advanced things

jimthompson5910 (jim_thompson5910):

well you got the basics down at least, so you can use that in some settings (like making your own desserts at home? that's useful I guess)

OpenStudy (anonymous):

can you factor 2x to the 4th power - 24x any other way besides 3x (x cubed - 8)

OpenStudy (anonymous):

ya what I learned will be useful

jimthompson5910 (jim_thompson5910):

2x^4 - 24x 2x(x^3 - 8) 2x(x^3 - 2^3) 2x(x - 2)(x^2 + 2x + 4) ... use the difference of cubes factoring rule

OpenStudy (anonymous):

*3x to the 4th power I mean

jimthompson5910 (jim_thompson5910):

yeah and you can decide to go back to it if you wanted to later down the road now that you have the basics under your belt

jimthompson5910 (jim_thompson5910):

oh 3x^4

jimthompson5910 (jim_thompson5910):

oh wait, lol made a serious typo above, ignore that...

OpenStudy (anonymous):

ok

jimthompson5910 (jim_thompson5910):

3x^4 - 24x 3x(x^3 - 8) 3x(x^3 - 2^3) 3x(x - 2)(x^2 + 2x + 4) ... use the difference of cubes factoring rule

jimthompson5910 (jim_thompson5910):

there's the correct version

OpenStudy (anonymous):

but how did you get the 2 cubed?

jimthompson5910 (jim_thompson5910):

8 = 2^3

jimthompson5910 (jim_thompson5910):

2^3 = 2*2*2 = 8

OpenStudy (anonymous):

oh right

OpenStudy (anonymous):

and if I wanted to solve for x, what would I do next?

jimthompson5910 (jim_thompson5910):

if 3x^4 - 24x = 0 then 3x(x - 2)(x^2 + 2x + 4) = 0

jimthompson5910 (jim_thompson5910):

3x(x - 2)(x^2 + 2x + 4) = 0 turns into 3x=0, or x-2=0 or x^2+2x+4=0

jimthompson5910 (jim_thompson5910):

you'll have to use the quadratic formula to solve x^2+2x+4=0 for x (you'll get 2 solutions here)

OpenStudy (anonymous):

and the quadratic formula is b = + - b times the square root of b squared - 4ac over 2a?

jimthompson5910 (jim_thompson5910):

close, the quadratic formula is \[\Large x = \frac{-b\pm\sqrt{b^2-4ac}}{2a}\]

OpenStudy (anonymous):

so it should be - 2 + - the square root of 2 squared - 4 times 1 times 4 over 2 x 1, which turns into -2 + - the square root of -12 over 2

OpenStudy (anonymous):

and now I have to find the square root of 12 which is is the square root of 4 times 3?

jimthompson5910 (jim_thompson5910):

good, you should get \[\Large x = \frac{-2\pm\sqrt{-12}}{2}\] and that turns into \[\Large x = \frac{-2+\sqrt{-12}}{2} \ \text{or} \ x = \frac{-2-\sqrt{-12}}{2}\] \[\Large x = \frac{-2+2i\sqrt{3}}{2} \ \text{or} \ x = \frac{-2-2i\sqrt{3}}{2}\] \[\Large x = -1+i\sqrt{3} \ \text{or} \ x = -1-i\sqrt{3}\]

jimthompson5910 (jim_thompson5910):

yes \[\large \sqrt{-12} = \sqrt{-1*4*3}\] \[\large \sqrt{-12} = \sqrt{-1}*\sqrt{4}*\sqrt{3}\] \[\large \sqrt{-12} = i*2*\sqrt{3}\] \[\large \sqrt{-12} = 2i*\sqrt{3}\]

OpenStudy (anonymous):

I finished all of my problems. Thanks so much for your help :)

jimthompson5910 (jim_thompson5910):

glad to be of help

OpenStudy (anonymous):

what did you major in, btw?

jimthompson5910 (jim_thompson5910):

math lol

jimthompson5910 (jim_thompson5910):

well math education really, learning to become a teacher

OpenStudy (anonymous):

would you want to teach high school or middle school?